Below is a table showing the daily production numbers for 1 worker in both Mexico and Canada. Use these numbers to answer the questions below.

1 worker in Mexico can produce in 1 day 1 worker in Canada can produce in 1 day
10 sodas and 5 pizzas 24 sodas and 8 pizzas

i. Determine a trade agreement for 1 pizza (price in terms of x number of sodas for 1 pizza) that both countries would benefit from. (Do not give me a range of prices. Just give me one price and one price only.) (Hint: who will be buying and who will be selling pizzas?)

j. Explain (using numbers) how Canada would benefit from your answer to (i) above.

k. Explain (using numbers) how Mexico would benefit from your answer to (i) above.
Answer #1

Opportunity cost of producing 1 pizza (Mexico) = 10/5 = 2 sodas

Opportunity cost of producing 1 pizza (Canada) = 24/8 = 3 sodas

i) A trade agreement where 1 pizza is exchanged for 2.5 sodas would benefit both the countries.

Mexico would be selling pizzas and Canada would be selling sodas.

j) Canada would benefit by = 3 - 2.5 = 0.5 sodas (for 1 pizza bought)

k) Mexico would benefit by = 2.5 - 2 = 0.5 sodas (for 1 pizza sold)